OpenStudy (anonymous):

Solve by completing the square (PLEASE show steps!!!) 9x^2+54x=18

OpenStudy (anonymous):

I would prefer steps and explanation, please

OpenStudy (anonymous):

9x^2+54x=18 first take 9 common from the L.H.S 9(x^2+6x)=18 9(x^2 + 2*x*3)=18 9(x^2 + 2*x*3 + 3^2 - 3^2) =18 (x+3)^2 -9 = 2

OpenStudy (anonymous):

Is it clear up to that?

OpenStudy (anonymous):

Yes, thank you

OpenStudy (anonymous):

So, u can continue from there?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

great